OpenStudy (caylie00):

solve the floolowing equation. u+2=4??

OpenStudy (anonymous):

u=2

OpenStudy (anonymous):

can you give me medal?

OpenStudy (caylie00):

could you explain it step by step

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

subtract 2 from both sides

OpenStudy (anonymous):

so..2+2=4

OpenStudy (anonymous):

and you get u=2

OpenStudy (anonymous):

2-2=0 and 4-2=2

OpenStudy (caylie00):

thank you so much